OpenStudy (anonymous):

Describe the transformation of the parent function f(x). h(x) = 0.25f(x)

hero (hero):

So f(x) is the parent function. And h(x) is equal to 1/4 of f(x). This means that whatever the value of h(x) for a given x, f(x) will always be 1/4 of h(x)'s value. For example, suppose f(x) = 16x Then h(x) = .25f(x) = .25(16x) = (1/4)(16x) = (16/4)x = 4x If you graph f(x) and h(x) you get https://www.desmos.com/calculator/zcum3bczvn Notice that when x = 1 f(x) = 16, h(x) = 4 When x = 2 f(x) = 32, h(x) = 8